OpenStudy (anonymous):

The length of a rectangle is three times its width. If the area of the rectangle is 75 yd^2, find its perimeter

OpenStudy (anonymous):

Kinda like the last one, L=3W.... so (3W)x(W)=75 to find W. Then Multiply Your W by 3 to find L.

OpenStudy (anonymous):

would it be 4w=75? no right??

OpenStudy (anonymous):

WxW is \[W ^{2}\]

OpenStudy (anonymous):

whats 3w^2 divided by 75..

OpenStudy (anonymous):

\[\frac{ 3W^2 }{ 3 }=\frac{ 75 }{ 3 }\]

OpenStudy (anonymous):

divide both sides by 3, so it is 75/3 = 25 which leaves you with \[W= \sqrt{25}\] so get rid of the exponent. \[W=5\]

OpenStudy (anonymous):

w=5?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

so L=15

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

then to find the perimiter do you still multiply L & W?

OpenStudy (anonymous):

and get 75

OpenStudy (anonymous):

So now do P=2L+2W, to find your perimeter.

OpenStudy (anonymous):

thats the area. how do i get perimiter

OpenStudy (anonymous):

p=40?

OpenStudy (anonymous):

Yea